OpenStudy (anonymous):

Find the slope of the line graphed below. A. 5/3 B. 3/5 C. -5/3 D. -3/5

Parth (parthkohli):

\( \color{Black}{\Rightarrow Slope = \Large{\Delta y \over \Delta x}}\)

Parth (parthkohli):

Simply find the change in \(y\) for every unit change in \(x\).

Parth (parthkohli):

Right. Can you just see my explanation and then find the slope?

mathslover (mathslover):

slope = tan\(\theta\) .. = \(\large{\frac{opposite}{adjacent}}\) calculare opposite by using pythagoras theorem ..

Parth (parthkohli):

Or you could just note down two points \((-3,0) \text{ and} (0,5)\)

mathslover (mathslover):

yes .. both r correct

mathslover (mathslover):

put opposite = 5 and adjacent = -3

mathslover (mathslover):

May i know what r u getting? @dasDaName ?

OpenStudy (lgbasallote):

@mathslover are you trying to find \(\theta\)?

Parth (parthkohli):

\[Slope = {y_2 - y_1 \over x _2 - x_1} \Longrightarrow {5 - 0 \over 0 - (-3) }\]

mathslover (mathslover):

no ..@lgbasallote .. yes @ParthKohli adjacent's place = 0-(-3)=3 ..

OpenStudy (lgbasallote):

then why propose tan theta?

Parth (parthkohli):

Why make it so complicated?

OpenStudy (anonymous):

haha, im not sure how to start this.do i just randomly pick to points on the graph thats on the line?

Parth (parthkohli):

Yes, the points that you can manage.

OpenStudy (lgbasallote):

weren't you always keen on going "by the book" and being very complicated @ParthKohli ? suddenly had a change of views? lol

Parth (parthkohli):

Two random points. Yep. It'd be better if the coordinates are integers.

OpenStudy (anonymous):

ok, (−3,0) and(0,5), then the next step would be? nah il try to manage it

Parth (parthkohli):

The slope formula is the next step.

Parth (parthkohli):

\[Slope = {y_2 - y_1 \over x_2 - x _1} \]

OpenStudy (anonymous):

\[(5-0)(0-(negative)3) \]

Parth (parthkohli):

Yes.

Parth (parthkohli):

There's a division sign in between, so careful.

OpenStudy (anonymous):

correct ?

Parth (parthkohli):

Yeah. What's 5 - 0?

OpenStudy (anonymous):

*5 and 3

OpenStudy (anonymous):

?

Parth (parthkohli):

Yeah

OpenStudy (anonymous):

alright, i really appreciate it man, thanks